OpenStudy (anonymous):

solve for x: sin3x=-cosx

OpenStudy (gorv):

it is sin^3 or just sin3x???

OpenStudy (anonymous):

sin3x..... Its not raised to the power of 3

OpenStudy (anonymous):

sin(3x) = sin(x+2x) this help?

OpenStudy (anonymous):

you have 2 identities to deal with

OpenStudy (anonymous):

@gopZ

OpenStudy (anonymous):

Cos(90-3x)= -cosx , because cos90-x=sinx ryt?

OpenStudy (dumbcow):

yes and -cos(x) = cos(180-x) so 90-3x = 180 - x solve for x

OpenStudy (dumbcow):

but that only captures 1 of 3 solutions

OpenStudy (anonymous):

Then lets find the general solution

OpenStudy (dumbcow):

sorry for delay...here use sum to product identity http://en.wikipedia.org/wiki/Trig_Identity#Product-to-sum_and_sum-to-product_identities \[\cos(90-3x) + \cos x = 0\] \[2\cos(\frac{90-2x}{2}) \cos(\frac{90-4x}{2}) = 0\] \[\cos(45-x) \cos(45-2x) = 0\] ...

OpenStudy (anonymous):

@dumbcow, ive been trying to award you with a medal but dis freakin' phone im using does not work properly with dis site! Anyway thanx for ur help

OpenStudy (dumbcow):

haha no worries :)